OpenStudy (anonymous):

write the equation of the line which has a slope of 3 and passes through ( -4,-7)

OpenStudy (anonymous):

(y+7)=3(x+4) y-3x=12-7 y-3x=5

OpenStudy (anonymous):

I need the answer in slope intercept form

OpenStudy (blurbendy):

y + 7 = 3(x +4) y + 7 = 3x + 12 y = 3x - 5 This is in slope-intercept form

OpenStudy (anonymous):

y=3x+5 not -5

OpenStudy (anonymous):

just rearrange it man

OpenStudy (blurbendy):

yes +5

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

Write the equation of the line passes through (6,-6)and(8,-5) slope intercept form